Inactive Tutor answered 06/28/20
Standardize the normal random variable.
z = (x - µ) / (σ / √n)
z = (11 - 10.8) / (3.4 / √20)
z = 0.26
Using our z-table, a z-score of 0.26 gives us a probability of .6026
The questions asks for the probability the mean life will be longer than 11 years. Which means we want the probability to the right of this z-score, so we need to subtract from 1.
1 - .6026 = .3974 or 39.74%